OpenStudy (anonymous):

Is the sequence geometric? 2, 6, 18, 54,... yes or no ?

OpenStudy (anonymous):

Yes..

OpenStudy (anonymous):

yes 2*3=6 6*3=18 18*3=54

OpenStudy (anonymous):

\[Common \; Ratio = \frac{6}{2} = \frac{18}{6} = \frac{54}{18} = 3\]